# Zero-Shot LLMs in Human-in-the-Loop RL: Replacing Human Feedback for Reward Shaping11footnotemark: 1

### v1

@@ Abstract
Reinforcement learning often faces challenges with reward misalignment, where agents optimize for given rewards but fail to exhibit the desired behaviors. This occurs when the reward function incentivizes proxy behaviors that diverge from the true objective. While human-in-the-loop (HIL) methods can help, they may exacerbate the problem, as humans are prone to biases that lead to inconsistent, subjective, or misaligned feedback, complicating the learning process. To address these issues, we propose two key contributions. First, we extend the use of zero-shot, off-the-shelf large language models (LLMs) for reward shaping beyond natural language processing (NLP) to continuous control tasks. By leveraging LLMs as direct feedback providers, we replace surrogate models trained on human feedback, which often suffer from the bias inherent in the feedback data it is trained on. Second, we introduce a hybrid framework ( LLM-HFBF) that enables LLMs to identify and correct biases in human feedback while incorporating this feedback into the reward shaping process. The LLM-HFBF framework creates a more balanced and reliable system by addressing both the limitations of LLMs (e.g., lack of domain-specific knowledge) and human supervision (e.g., inherent biases). By enabling human feedback bias flagging and correction, our approach improves reinforcement learning performance and reduces reliance on potentially biased human guidance. Empirical experiments show that biased human feedback significantly reduces performance, with average episodic reward (AER) dropping from 28.472 in (unbiased approaches) to 7.039 (biased with conservative bias). In contrast, LLM-based approaches maintain a matching AER like unbiased feedback, even in custom edge case scenarios.
@@ 1 Introduction
Figure 1: Performance comparison of direct human feedback (HF-D), biased human feedback (Biased HF-D(AGG) and Biased HF-D(CON)), and proposed LLM-based feedback methods (LLM-D and LLM-HFBF) for reward shaping and RL training.
Personal human biases and preferences can distort feedback, preventing optimal learning [5]. Several approaches have tried to train and use surrogate models as a proxy to direct human feedback to make HIL RL more cost effective, for example Reinforcement Learning from AI Feedback (RLAIF) [7] and Iterative Reward Shaping (ITERS) [8]. However, these methods face significant limitations: ITERS follows human feedback without accounting for potential biases, while RLAIF relies entirely on LLM-generated feedback which may lack the nuanced human understanding required in certain complex scenarios. Both approaches can be compromised when the feedback itself suffers from human bias, as the surrogate models inevitably inherit and potentially amplify these biases during training.
In this work, we first demonstrate that off-the-shelf LLMs can effectively shape rewards in continuous control tasks within the MuJoCo environment, requiring no fine-tuning (zero-shot). However, recognizing that human feedback remains essential in certain scenarios—particularly in high-risk and sensitive environments—we introduce a hybrid framework where LLMs identify potential biases in human feedback, enabling targeted bias correction. To illustrate the impact of different feedback strategies, Figure 1 compares the performance of direct human feedback (HF-D), human feedback influenced by aggressive and conservative biases (Biased HF-D(AGG) and Biased HF-D(CON), respectively), and our proposed LLM-based methods: a direct reward shaping approach (LLM-D) and a hybrid feedback framework that flags and corrects biases (LLM-HFBF).
@@
    We validate our approach in continuous control tasks (highway navigation in MuJoCo), demonstrating that zero-shot LLMs can effectively shape rewards in these domains at par their unbiased counterparts.
@@ 4 Methodology
To investigate bias-induced performance degradation in feedback systems, particularly within the context of reward shaping, and to assess their application with large language models (LLMs), we implement three distinct behavioral profiles using rule-based proxies: IDEAL (balanced), AGGRESSIVE (risk-seeking), and CONSERVATIVE (safety-prioritizing). Each profile is characterized by style coefficients that influence lane changes, collision avoidance, and speed optimization. We compare four feedback integration strategies: HF-RSM (human feedback with learned reward shaping model), HF-D (direct human feedback), LLM-D (direct LLM feedback), and LLM-HFBF (hybrid approach where LLMs detect and adapt to human feedback biases). Additionally, we design custom edge case scenarios to evaluate how these strategies handle situations where rule-based systems struggle with contextual adaptation.
@@ 4.1.3 Human Feedback - Direct (HF-D)
Here the rule based proxy is used directly without the surrogate model. Though rule-based systems may not be ideal for most cases, they are well suited for our purpose, where we address the challenges of using direct human feedback in reward shaping scenarios, particularly the risk of bias in such feedback.
@@
In the Highway environment (see Sec. 5.1), drawing motivation from the rules used in [8] we have used the following to provide feedback as a stand-in for human input, i.e. three behavioral profiles, including the ideal (default) rule and two biased behaviors. They are: IDEAL (default), AGGRESSIVE or AGG-encourages risky behaviors, and CONSERVATIVE or CON- prioritizes safety and avoids risks. AGG and CON represents biased feedback profiles, based on their specific driving styles, the style coefficients $B$ are modified, thus affecting the resulting $\zeta_{\text{HF-D}}$.
@@
where $p\in\{\text{IDEAL},\text{AGG},\text{CON}\}$ is the parameter set, and $B_{\text{lane}}^{p}$, $B_{\text{collision}}^{p}$, and $B_{\text{speed}}^{p}$ are the style coefficient arrays for parameter set $p$. Moreover $w_{lane}$, $w_{collision}$, and $w_{speed}$ are weights (default=1).
Note on Scenario Indices and Parameters: The table indices map to scenarios defined by several parameters: time-to-collision ($ttc$), acceleration ($acc$), lane change flag ($lc$), traffic density ($\rho$), agent speed ($v$), and speed thresholds ($v_{thresh}$). For Lane Changes, $n_{0}$-$n_{3}$ represent 0-3 lane changes. For Collision Avoidance, $c_{0}$-$c_{4}$ map to: potential collision ($0.5\leq ttc\leq 2.0$) with $acc<0$ ($c_{0}$), $acc>0$ ($c_{1}$), or $lc=1$ ($c_{2}$); immediate risk ($ttc<0.5$, $c_{3}$); and safe path ($ttc>2.0$, $c_{4}$). For Speed Optimization, $s_{0}$-$s_{8}$ combine traffic density levels ($\rho\in\{1,2,3\}$) with speed categories relative to thresholds: $v>v_{thresh}[1]$ (high), $v_{thresh}[0]\leq v\leq v_{thresh}[1]$ (medium), and $v<v_{thresh}[0]$ (low).
@@
$IDEAL$ +2 0 -1 -2 +1 0 +1 -2 +2 +2 +1 -1 -1 +2 -1 -2 -1 +2
$AGG$2 +1 +2 +2 -2 +2 +1 -1 -2 +2 +1 -2 +1 0 -1 +2 +1 -2
$CON$2 +1 +2 +2 -2 +2 +1 -1 +2 +2 +1 -2 +1 0 -1 0 +2 +1
This table presents the style coefficient values across three different driving profiles: IDEAL, AGG, and CON. The matrix is organized into three behavioral categories: Lane Change behavior ($B_{\text{lane}}^{p}$, based on number of lane changes $n$), Collision Avoidance strategies ($B_{\text{collision}}^{p}$, based on time-to-collision $t_{\text{tc}}$, acceleration $a_{\text{cc}}$, and lane change flag $l_{\text{c}}$), and Speed Optimization ($B_{\text{speed}}^{p}$, based on traffic density $\rho$ and agent speed $v$ relative to threshold speeds $v_{\text{thresh}}$). Each entry represents the style coefficient value for a specific scenario, with positive values encouraging behaviors and negative values discouraging them.
@@
    Conservative Feedback: Promotes moderate lane changes, rewards safe paths, and discourages excessive speeding in heavy traffic or slowing down in collisions. It aims for a balance between speed and safety, encouraging efficient driving that conserves fuel while avoiding unnecessary risks. The efficient feedback is encoded through coefficients like rewarding moderate lane changes and safe paths (e.g., positive values for $B_{\text{lane}}[n_{1},n_{2}]$) while penalizing extreme behaviors like excessive speed in high-traffic conditions (e.g., adjusting $B_{\text{speed}}[s_{0},s_{1}]$).
@@ 2. Reward Appropriateness Evaluation:
Expected Format:
@@
Analysis: The action taken by the agent at this time step (5) was to do nothing (Action = 0). Since no explicit control action is specified in this dataset, we infer its effect based on the context. A collision occurred (Collision Flag = 1), so it’s reasonable to assume that the action was ineffective in avoiding the collision. The agent could have changed lanes or adjusted speed to prevent the crash.
@@ 4.1.4 Large Language Model - Direct (LLM-D)
In this case, we use a LLM (off-the-shelf ) to provide feedback for reward shaping. We do not finetune or train the LLM in any way and use it directly, integrating with our framework via prompts. The LLM acts as a proxy for direct human feedback, offering feedback on agent actions and rewards based on structured trajectory data. The LLM receives a data frame where each row contains the episode number, time step, state, action, reward, next state, collision flag, and lane index. The LLM generates feedback based on the state-action pair at each time step, evaluating action effectiveness and reward appropriateness. It categorizes actions as effective, ineffective, or neutral, with adjustments of $+0.5$, $-0.5$, or 00, and assesses reward appropriateness (+0.2 or -0.2).
@@
Your Task:
@@
Sample Output
The Adjusted score (0.0) does not align with the agent’s impact on trajectory, as indicated by the principal components (PC1, PC2, PC3). PC1 and PC3 suggest some level of speed and trajectory consistency, while PC2 indicates smooth driving but with room for improvement.
Considerations:
1. Reward (0.7333) indicates the action was somewhat beneficial.
2. Collision Flag (0) shows no collision, indicating safe driving.
3. Lane Index (3) suggests the agent stayed in lane 3.
4. PC1 and PC3 reflect some deviation from ideal values, suggesting room for improvement.
5. PC2 reflects some reward variation, which could be improved with smoother actions.
@@
llm_score = 0.3 (encouraging improvement)
Table 3: Examples on how LLM in the LLM-HFBF approach evaluates the input data and adjusts the reward score. The input includes details such as episode number, time step, state, action, reward, collision flag, and compressed representation in form of principal components (PC1, PC2, PC3). In the output, the LLM identifies that the provided adjusted score (0.0) does not align with the agent’s actions as reflected in the principal components. It flags the score as biased and adjusts it to 0.3, suggesting a more appropriate score based on reinforcement learning principles. This adjusted score encourages continued learning while acknowledging areas for improvement
@@ 4.1.5 LLM Aided Human Feedback Bias Flagging (LLM-HFBF)
The LLM-HFBF approach represents our hybrid framework that leverages LLMs to detect and correct potential biases in human feedback (see Table 3). This approach can be formalized as follows: Let
@@ 5.1 Experiment Setup
The experiments were designed to evaluate the efficacy of LLM-driven reward shaping in comparison to traditional human feedback-driven reward shaping across multiple RL environments. Here we describe the different scenarios we set the environment to test the efficacy of the different feedback mechanisms:
@@ A.1. Highway (Highway-v0)
A.1. Default environment The default Highway environment is a multi-lane driving scenario with four lanes and a total of 50 vehicles, where the agent controls one vehicle. The agent’s initial lane ID is unspecified, allowing flexible starting positions. The reward system includes a collision penalty of -1, no direct reward for lane changes (0 reward), and an incentive for staying in the right lane with a 0.1 reward. High-speed driving is encouraged with a reward of 0.4 for maintaining speeds within the 20 to 30 m/s range. Other vehicles follow the Intelligent Driver Model (IDM) for realistic traffic behavior.

### v2

@@ Abstract
Reinforcement learning (RL) often struggles with reward misalignment, where agents optimize given rewards but fail to exhibit the desired behaviors. This arises when the reward function incentivizes proxy behaviors misaligned with the true objective. While human-in-the-loop (HITL) methods can mitigate this issue, they also introduce biases, leading to inconsistent and subjective feedback that complicates learning. To address these challenges, we propose two key contributions. First, we extend the use of zero-shot, off-the-shelf large language models (LLMs) for reward shaping beyond natural language processing (NLP) to continuous control tasks. Using LLMs as direct feedback providers eliminates the need for surrogate models trained on human feedback, which often inherit biases from training data. Second, we introduce a hybrid framework (LLM-HFBF) that enables LLMs to identify and correct biases in human feedback while incorporating this feedback into the reward shaping process. The LLM-HFBF framework creates a more balanced and reliable system by addressing both the limitations of LLMs (e.g., lack of domain-specific knowledge) and human supervision (e.g., inherent biases). By enabling human feedback bias flagging and correction, our approach improves reinforcement learning performance and reduces reliance on potentially biased human feedback.
Empirical experiments show that biased human feedback significantly reduces performance, with Average Episodic Reward dropping by nearly 94% compared to unbiased approaches. In contrast, LLM-based methods sustain performance at a similar level to unbiased feedback, even in challenging edge-case scenarios.
@@ 1 Introduction
Figure 1: Performance comparison of direct human feedback (HF-D), biased human feedback (Biased HF-D(AGG) and Biased HF-D(RAD)), and proposed LLM-based feedback methods (LLM-D and LLM-HFBF) for reward shaping and RL training. Results are from Highway-v0 environment.
Personal human biases and preferences can distort feedback, preventing optimal learning [5]. Several approaches have tried to train and use surrogate models as a proxy to direct human feedback to make HITL-RL more cost effective, for example Reinforcement Learning from AI Feedback (RLAIF) [7] and Iterative Reward Shaping (ITERS) [8]. However, these methods face significant limitations: ITERS follows human feedback without accounting for potential biases, while RLAIF relies entirely on LLM-generated feedback which may lack the nuanced human understanding required in certain complex scenarios. Both approaches can be compromised when the feedback itself suffers from human bias, as the surrogate models inevitably inherit and potentially amplify these biases during training.
While LLMs can inherit biases from their pretraining corpora [9], including behavioral and social biases such as ingroup favoritism [10] and systematic biases in opinion formation exploitable through adversarial attacks [11], they also offer distinct advantages over human feedback in reinforcement learning contexts. Human feedback is often inconsistent, subjective, and variable over time, whereas LLM biases tend to be systematic and therefore potentially auditable. Although alignment techniques such as instruction tuning and preference modeling do not fully eliminate LLM biases [10], they establish a foundation for generating feedback that is more predictable and scalable.
In this work, we first demonstrate that off-the-shelf LLMs can effectively shape rewards in continuous control tasks within the MuJoCo environment, requiring no fine-tuning (zero-shot). However, recognizing that human feedback remains essential in certain scenarios—particularly in high-risk and sensitive environments—we introduce a hybrid framework where LLMs identify potential biases in human feedback, enabling targeted bias correction. To illustrate the impact of different feedback strategies, Figure 1 compares the performance of direct human feedback (HF-D), human feedback influenced by aggressive and reckless adaptive biases (Biased HF-D(AGG) and Biased HF-D(RAD), respectively), and our proposed LLM-based methods: a direct reward shaping approach (LLM-D) and a hybrid feedback framework that flags and corrects biases (LLM-HFBF). The code associated with this paper is publicly available [^2] https://github.com/RizanSM/zero_shot_llms_in_HIL_RL.
@@
    We validate our approach in continuous control tasks (Highway-v0 and Reacher-v5 in MuJoCo), demonstrating that zero-shot LLMs can effectively shape rewards in these domains at par their unbiased counterparts.
@@ 4 Methodology
To investigate bias-induced performance degradation in feedback systems, particularly within the context of reward shaping, and to assess their application with large language models (LLMs), we implement three distinct behavioral profiles using rule-based proxies: IDEAL (balanced), AGGRESSIVE (risk-seeking), and RECKLESS ADAPTIVE (impulsive opportunistic). Each profile is characterized by style coefficients that influence lane changes, collision avoidance, and speed optimization. We compare four feedback integration strategies: HF-RSM (human feedback with learned reward shaping model), HF-D (direct human feedback), LLM-D (direct LLM feedback), and LLM-HFBF (hybrid approach where LLMs detect and adapt to human feedback biases). Additionally, we design custom edge case scenarios to evaluate how these strategies handle situations where rule-based systems struggle with contextual adaptation.
@@ 4.1.3 Human Feedback - Direct (HF-D)
Here the rule-based proxy is used directly without the surrogate model. Although rule-based systems may not be ideal for most cases, they are well suited for our purpose, where we address the challenges of using direct human feedback in reward shaping scenarios, particularly the risk of bias in such feedback.
@@
In the Highway environment (see Sec. 5.1), drawing motivation from the rules in [8], we implemented three behavioral profiles to simulate human feedback: IDEAL (default), AGGRESSIVE (AGG), and RECKLESS ADAPTIVE (RAD). The AGG and RAD profiles serve as biased feedback proxies, with AGG encouraging risk-seeking behaviors such as rapid lane changes and higher speeds, and RAD promoting impulsive and opportunistic driving through abrupt maneuvers and inconsistent speed choices. These profiles are empirically grounded in patterns of risky and erratic driving identified in [30], which link psychological traits to driving behaviors. Each biased profile is computationally simulated by modifying the style coefficients $B$, thereby affecting the resulting $\zeta_{\text{HF-D}}$.
@@
where $p\in\{\text{IDEAL},\text{AGG},\text{RAD}\}$ is the parameter set, and $B_{\text{lane}}^{p}$, $B_{\text{collision}}^{p}$, and $B_{\text{speed}}^{p}$ are the style coefficient arrays for parameter set $p$. Moreover $w_{lane}$, $w_{collision}$, and $w_{speed}$ are weights (default=1).
Note on Scenario Indices and Parameters: The table indices map to scenarios defined by several parameters: time-to-collision (ttcttc), acceleration (accacc), lane change flag ($lc$), traffic density ($\rho$), agent speed ($v$), and speed thresholds ($v_{thresh}$). For Lane Changes, $n_{0}$-$n_{3}$ represent 0-3 lane changes. For Collision Avoidance, $c_{0}$-$c_{4}$ map to: potential collision ($0.5\leq ttc\leq 2.0$) with $acc<0$ ($c_{0}$), $acc>0$ ($c_{1}$), or $lc=1$ ($c_{2}$); immediate risk ($ttc<0.5$, $c_{3}$); and safe path ($ttc>2.0$, $c_{4}$). For Speed Optimization, $s_{0}$-$s_{8}$ combine traffic density levels ($\rho\in\{1,2,3\}$) with speed categories relative to thresholds: $v>v_{thresh}[1]$ (high), $v_{thresh}[0]\leq v\leq v_{thresh}[1]$ (medium), and $v<v_{thresh}[0]$ (low).
@@
IDEALIDEAL +1 0 -1 -1 +1 0 +1 -2 +2 +2 +1 -1 -1 +2 -1 -2 -1 +2
AGGAGG -2 +1 +2 +2 -2 +2 +1 -1 -2 +2 +1 -2 +1 0 -1 +2 +1 -2
RADRAD -2 0 +1 +1 -2 +1 +1 0 -2 +1 0 -2 0 -1 -2 +1 0 -2
This table presents the style coefficient values across three different driving profiles: IDEAL, AGG, and RAD. The matrix is organized into three behavioral categories: Lane Change behavior ($B_{\text{lane}}^{p}$, based on number of lane changes $n$), Collision Avoidance strategies ($B_{\text{collision}}^{p}$, based on time-to-collision $t_{\text{tc}}$, acceleration $a_{\text{cc}}$, and lane change flag $l_{\text{c}}$), and Speed Optimization ($B_{\text{speed}}^{p}$, based on traffic density $\rho$ and agent speed $v$ relative to threshold speeds $v_{\text{thresh}}$). Each entry represents the style coefficient value for a specific scenario, with positive values encouraging behaviors and negative values discouraging them.
@@
    Reckless Adaptive Feedback (RAD): Promotes impulsive and opportunistic driving by dynamically rewarding abrupt maneuvers and inconsistent speed choices in varying traffic conditions. The strategy prioritizes short-term gains with minimal regard for potential hazards. Safe behaviors, such as cautious path selection or slowing down to avoid collisions, are penalized, while risky actions—including sudden accelerations and frequent multi-lane changes—are rewarded regardless of safety. This behavior is implemented by adjusting the style coefficients $B$: coefficients corresponding to safe actions (e.g., $B_{\text{collision}}[c_{0},c_{4}]$ for cautious path selection and slowing down) are assigned negative values, whereas coefficients associated with abrupt, high-risk maneuvers are positively reinforced.
@@ 2. Reward Appropriateness Evaluation:
Expected Format: JSON object containing the following:
@@
Analysis: The action taken by the agent in the fourth lane (Lane Index = 3) led to a collision (Collision flag = 1). Since no explicit control action is specified in this dataset, we infer its effect based on the context. It’s reasonable to assume that the action was ineffective in avoiding the collision. The agent could have changed lanes or adjusted speed to prevent the crash.
@@ 4.1.4 Large Language Model - Direct (LLM-D)
In this case, we use a LLM (off-the-shelf ) to provide feedback for reward shaping. We do not finetune or train the LLM in any way and use it directly, integrating with our framework via prompts. The LLM acts as a proxy for direct human feedback, offering feedback on agent actions and rewards based on structured trajectory data. The LLM receives a data frame where each row contains the episode number, time step, state, action, reward, next state, collision flag, and lane index. The LLM generates feedback based on the state-action pair at each time step, evaluating action effectiveness and reward appropriateness. It categorizes actions as effective, ineffective, or neutral, with adjustments of +0.5+0.5, 0.5-0.5, or 0, and assesses reward appropriateness (+0.2 or -0.2).
@@
Your Task: JSON objects of the following:
- VERDICT
@@
LLM $\rightarrow$ Output
VERDICT: BIASED SCORE ALLOTTED
Justification: The presence of a Collision Flag indicates an undesirable event, which has not been penalized in the Adjusted Score (4.0). Although the reward received (0.066) suggests that the agent’s action yielded a short-term gain, the llm score(-1.5) penalizes it and emphasizes safety over immediate outcomes. The principal components reinforce this view: PC3 highlights poor trajectory-level consistency, whereas PC1 reflects acceptable control over speed and lane changes.
@@
llm_score = -1.5
Considerations:
1. Adjusted Score reflects the human feedback in the form of a composite value that integrates lane changes, collision avoidance, and speed optimization.
2. Reward (0.066) is positive, showing-term beneficial outcome .
3. PC1 suggests acceptable control over speed and lane changes.
4. PC2 reflects moderate stability of reward variation across steps.
5. PC3 highlights poor trajectory-level consistency, alligning with the observed collision and framented beh
Table 3: Examples on how LLM in the LLM-HFBF approach evaluates the input data and adjusts the reward score. The input includes details such as episode number, time step, state, action, reward, collision flag, and compressed representation in form of principal components (PC1, PC2, PC3). In the output, the LLM identifies that the provided adjusted score (4.0) does not align with the agent’s actions as reflected in the principal components. It flags the score as biased and adjusts it to -1.5, suggesting a more appropriate score based on reinforcement learning principles. This adjusted score encourages continued learning while acknowledging areas for improvement.
@@ 4.1.5 LLM Aided Human Feedback Bias Flagging (LLM-HFBF)
The LLM-HFBF approach represents our hybrid framework that leverages LLMs to detect and correct potential biases in human feedback (see Fig. 2 for the LLM-HFBF workflow and Table 3 shows the prompt and corresponding LLM feedback). This approach can be formalized as follows: Let
@@
The objective representation of trajectory-level agent behavior in the LLM-HFBF approach is derived from the following raw features: (1) State, (2) Action, (3) Reward, (4) Collision flag, (5) Lane index, and (6) Adjusted rewards (computed via biased reward shaping). To extract the most informative structure from this high-dimensional feature space, Principal Component Analysis (PCA) was applied. The first three principal components (PCs) capture 95.17% of the total variance, with PC1 accounting for 45.39%, PC2 for 27.92%, and PC3 for 21.86%. This dimensionality reduction facilitates efficient clustering and analysis by preserving essential behavioral patterns—including speed regulation, lane-change strategies, collision risk, and overall trajectory consistency—while filtering out redundant correlations across features.
@@ 5.1 Experiment Setup
Below we describe the different RL environments, custom edge case scenarios, and metrics we have used to evaluate and compare the performance of the different FIS.
@@ A.1. Highway (Highway-v0)
A.1. Default environment: It is a multi-lane driving scenario with four lanes and a total of 50 vehicles, where the agent controls one vehicle. The agent’s initial lane ID is unspecified, allowing flexible starting positions. The reward function includes a collision penalty of -1, no direct reward for lane changes (0 reward), and an incentive for staying in the right lane with a 0.1 reward, given by
[…]
High-speed driving is encouraged with a reward of 0.4 for maintaining speeds within the 20 to 30 m/s range. Other vehicles follow the Intelligent Driver Model (IDM) for realistic traffic behavior.